# Turing Video-based Cognitive Tests to Handle Entangled Concepts

### v1

@@ 2 Identifying entanglement in physics and cognition
At an experimental level, we performed cognitive tests on human participants [30, 31, 38, 39], document retrieval tests on structured corpuses of documents [35, 37] and image retrieval tests on web search engines [33, 36, 39] using various combinations of two concepts. We mainly investigated the conceptual combination The Animal Acts, which we considered as a composite conceptual entity made up of the individual conceptual entities Animal and Acts. The tests had the form of the Bell-type test above, and we found (i) a systematic violation of the CHSH inequality in Equation (1). However, we also found (ii) a systematic violation of the marginal law (Equations (3) and (4)) and, in some cases, (iii) the CHSH factor in Equation (2) exceeded Cirel’son’s bound.
@@
While empirical finding (i) substantially agreed with the predictions of quantum theory and indicated the presence of ‘conceptual entanglement’, empirical findings (ii) and (iii) were unexpected, as they are not believed to occur in quantum physics. This led us to initiate a theoretical investigation on a problem that is usually overlooked in physics, the ‘identification problem’, that is, the problem of recognising individual entities of a composite entity by performing on the latter the typical coincidence measurements of Bell-type tests [31]. We thus elaborated a general theoretical framework to model any Bell-type situation, independently of the nature, physical or conceptual, of the entities involved, within the mathematical formalism of quantum theory in Hilbert space [31, 34]. In this theoretical framework, one applies the standard prescription of quantum theory according to which the composite entity $S_{12}$ needs to be associated with a complex Hilbert space whose dimension is determined by the number of distinct outcomes of the performed measurements. Since in principle, each of the coincidence measurements $AB$, $AB^{\prime}$, $A^{\prime}B$ and $A^{\prime}B^{\prime}$ have four distinct outcomes, $S_{12}$ should be represented in the Hilbert space $\mathbb{C}^{4}$ of all ordered 4-tuples of complex numbers. Only in the attempt of ‘recognizing’ individual entities $S_{1}$ and $S_{2}$ within the composite entity, one considers possible isomorphisms with the tensor product Hilbert space $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, where each copy of $\mathbb{C}^{2}$ takes into account the fact that measurements with two distinct outcomes can be performed on $S_{1}$ and $S_{2}$ in a Bell-type setting. And it is only at this stage, i.e. when individual entities are recognized from measurements performed on the composite entity, that entanglement can be identified. We proved in [31], that no unique isomorphism exists between $\mathbb{C}^{4}$ and $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, and this is the reason why, from a mathematical point of view, different ways exist to account for entanglement being present within the composite entity $S_{12}$ with respect to the individual entities $S_{1}$ and $S_{2}$ that are recognized as parts of $S_{12}$.
@@
Essentially, entanglement manifests itself when the probabilities of a coincidence measurement on $S_{12}$ cannot be written as products of probabilities of measurements on the component entities $S_{1}$ and $S_{2}$. Hence, entanglement is a property of the relation between the coincidence measurements and measurements on the component entities. Only when the additional symmetry mentioned above, i.e. the marginal law being satisfied, is present in all coincidence measurements, the entanglement of these different coincidence measurements can be captured in a state of the composite entity. If this is true for all coincidence measurements, one can prove that there is only one isomorphism connecting $\mathbb{C}^{4}$ with $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, and $\mathbb{C}^{4}$ can be directly replaced by $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ in that case. This means that the situation usually reported in modern manuals of quantum theory is exceptional and not the general one [31]. In this general situation, where the marginal law is empirically violated, no unique isomorphism exists between $\mathbb{C}^{4}$ and $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$.
@@
The quantum theoretical framework above enables modelling of empirical data collected in Bell-type tests where the marginal law does not hold and the violation of the CHSH inequality in Equation (1) exceeds Cirel’son’s bound, in those tests where this occurs, by introducing and representing entangled measurements. In particular, we have proved that, whenever the concepts Animal and Acts combine to form the combination The Animal Acts, a strong form of entanglement is created between Animal and Acts, which is such that, not only the state of the composite entity The Animal Acts is entangled, but also the coincidence measurements are entangled [31, 36, 37, 38].
@@
We will see in Sections 4 and 5 that the entanglement above is due to the peculiar way in which the meaning of The Animal Acts relates to the meanings of Animal and Acts, which violates the classical semantic rules of composition. Before doing this, however, we need to present the details of the novel cognitive test based on videos which we have recently performed on this conceptual combination. This will be the aim of Section 3.
@@ 3 A novel video-based cognitive test
A sample of 221 individuals were presented in a ‘within subjects design’ a HTML5 questionnaire which contained four coincidence experiments $AB$, $AB^{\prime}$, $A^{\prime}B$ and $A^{\prime}B^{\prime}$ whose setting was similar to the typical setting of a Bell-type test presented in Section 2. More specifically, participants were preliminarily asked to read an ‘introductory text’ where an explanation of the type of judgement test they had to complete and a description of the tasks involved in the judgement test were provided. More specifically, the participants had to preliminary complete a simple introductory test on the concept Fruit, judging the item that they considered as a good example of the concept Fruit. The items of Fruit, indicated in Figure 1, were generated by image-based generative intelligence, and we proposed the item Strawberry, which is an enlarged portion of an inflorescence, and Tomato, which is a fruit, to draw maximum attention to the choice. Cherry and Pomegranate were the remaining items of Fruit to judge upon.
@@
Then, in each coincidence measurement, participants were asked to choose which item in a list of four items they judged as a good example of the conceptual combination The Animal Acts. In each coincidence measurement, each of the four possible items of The Animal Act was a video showing an animal and the corresponding action. More specifically, the videos, indicated in Figure 2, were created as follows:
@@
Figure -2: sub-test 1
Figure -1: sub-test 2
@@
Before analysing the results of the test, it is worth to make some remarks on the choice of its setting. The detection of entanglement in Bell-type cognitive tests, indeed, relies on a precise preparation of both the initial state and the coincidence measurements. Insufficient initial conditions, as lacking a brief introduction and prior knowledge of the animals and sounds involved, can hinder the ability to identify entanglement. Birds could offer an interesting case study because some species possess unique vocalizations distinct from other individuals. This individuality makes them well-suited for entanglement experiments involving experts. One could, e.g., consider, two pairs of items of Animal, namely, (Cuckoo, Nightingale) and (Thrush, Goldfinch), and two pairs of items of Acts, (Cuckoos, Trills) and (Whistles, Snails), without determining a violation of the CHSH inequality in Equation (1). In other words, cognitive tests on unskilled participants may fail to detect entanglement or yield lower value of the CHSH factor $\Delta_{CHSH}$ in Equation (2). A similar situation is likely to occur if the actions are not typical of the animals. Consider, e.g., two pairs of items of Animal (Horse, Bear) and (Dog, Camel), and two pairs of items of Sport (Racing, Fishing) and (Sledding, Show Jumping) for the conceptual combination The Animal does Sport. Also in this case, we expect the CHSH inequality in Equation (1) not to be violated or to show a little violation. This is why the identification of the items to use in the cognitive test followed strict prescriptions, as follows:
@@
In the coincidence measurement $AB$, participants had to choose the best example of The Animal Acts within the four items:
@@
In the coincidence measurement $AB^{\prime}$, participants had to choose the best example of The Animal Acts within the four items:
@@
In the coincidence measurement $A^{\prime}B$, participants had to choose the best example of The Animal Acts within the four items:
@@
Finally, in the coincidence experiment $A^{\prime}B^{\prime}$, participants had to choose the best example of The Animal Acts within the four items:
@@
For each coincidence measurement $XY$, we collected the relative frequencies of the obtained responses which we considered, in the large number limit, as the probability $\mu(X_{i}Y_{j})$ that the outcome $X_{i}Y_{j}=\pm 1$ is obtained in the corresponding measurement, $X=A,A^{\prime}$, $Y=B,B^{\prime}$. Table 1 reports the judgement probabilities computed in this way. Referring to these probabilities, we can then calculate the expectation values, or correlation functions, of the coincidence measurements $AB$, $AB^{\prime}$, $A^{\prime}B$ and $A^{\prime}B^{\prime}$, as follows:
@@
The numerical value $2.9774$ exceeds the classical limit imposed by the CHSH inequality in Equation (1) and is also above Cirel’son’s bound $2\sqrt{2}\approx 2.8284$. A simple check in Table 1 reveals that the marginal law in Equations (3) and (4) is also systematically violated here. This empirical pattern confirms and strengthens the results obtained in previous text-based cognitive tests on the conceptual combination The Animal Acts, namely [30, 38, 39]. Table 2 reports the judgement probabilities and the CHSH factor in these three text-based tests. By comparing the CHSH factors in Tables 1 and 2, we can see that a systematic violation of the CHSH inequality occurs in all tests. We also notice that the value of the CHSH factor in the test in [30] is far from Cirel’son’s bound, the test in [38] is close to that bound, and the test in [39], together with the present test violate Cirel’son’s bound. That the test in [30] has a relatively lower deviation from classicality than the other tests might be due to the fact that the introductory text in [30] encouraged the participants to also take into account emotions and imagination in their judgements. This led them to make choices that are less natural, as The Bear Meows, The Cat Growls, or The Tiger Meows, thus determining a relatively lower violation of the CHSH inequality in Equation (1).

### v2

@@ 2 Identifying entanglement in physics and cognition
At an experimental level, we performed cognitive tests on human participants (Aerts & Sozzo,, 2011, 2014; Aerts et al., 2023b, ; Bertini Lemporini & Moriani,, 2023), document retrieval tests on structured corpuses of documents (Beltran & Geriente,, 2019; Aerts et al.,, 2021) and image retrieval tests on web search engines (Aerts Arguëlles,, 2018; Aerts Arguëlles & Sozzo,, 2020; Bertini Lemporini & Moriani,, 2023) using various combinations of two concepts. We mainly investigated the conceptual combination The Animal Acts, which we considered as a composite conceptual entity made up of the individual conceptual entities Animal and Acts. The tests had the form of the Bell-type test above, and we found (i) a systematic violation of the CHSH inequality in Equation (1). However, we also found (ii) a systematic violation of the marginal law conditions in Equations (3) and (4) and, in some cases, (iii) the CHSH factor in Equation (2) exceeded Cirel’son’s bound.
@@
While empirical finding (i) substantially agreed with the predictions of quantum theory and indicated the presence of ‘conceptual entanglement’, empirical findings (ii) and (iii) were unexpected, as they are not believed to occur in quantum physics. This led us to initiate a theoretical investigation on a problem that is usually overlooked in physics, the ‘identification problem’, that is, the problem of recognising individual entities of a composite entity by performing on the latter the typical joint measurements of Bell-type tests (Aerts & Sozzo,, 2014). We thus elaborated a general theoretical framework to model any Bell-type situation, independently of the nature, physical or cognitive, of the entities involved, within the mathematical formalism of quantum theory in Hilbert space (Aerts & Sozzo,, 2014; Aerts et al.,, 2019). In this theoretical framework, one applies the standard prescription of quantum theory according to which the composite entity $S_{12}$ needs to be associated with a complex Hilbert space whose dimension is determined by the number of distinct outcomes of the performed measurements. Since in principle, each of the joint measurements $AB$, $AB^{\prime}$, $A^{\prime}B$ and $A^{\prime}B^{\prime}$ have four distinct outcomes, $S_{12}$ should be represented in the Hilbert space $\mathbb{C}^{4}$ of all ordered 4-tuples of complex numbers. Only in the attempt of ‘recognizing’ individual entities $S_{1}$ and $S_{2}$ within the composite entity, one considers possible isomorphisms with the tensor product Hilbert space $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, where each copy of $\mathbb{C}^{2}$ takes into account the fact that measurements with two distinct outcomes can be performed on $S_{1}$ and $S_{2}$ in a Bell-type setting. And it is only at this stage, i.e. when individual entities are recognized from measurements performed on the composite entity, that entanglement can be identified. We proved in Aerts & Sozzo, (2014), that no unique isomorphism exists between $\mathbb{C}^{4}$ and $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, and this is the reason why, from a mathematical point of view, different ways exist to account for entanglement being present within the composite entity $S_{12}$ with respect to the individual entities $S_{1}$ and $S_{2}$ that are recognized as parts of $S_{12}$.
@@
Essentially, entanglement manifests itself when the probabilities of a joint measurement on $S_{12}$ cannot be written as products of probabilities of measurements on the component entities $S_{1}$ and $S_{2}$. Hence, entanglement is a property of the relation between the joint measurements and measurements on the component entities. Only when the additional symmetry mentioned above, i.e. the marginal law conditions being satisfied, is present in all joint measurements, the entanglement of these different joint measurements can be captured in a state of the composite entity. If this is true for all joint measurements, one can prove that there is only one isomorphism connecting $\mathbb{C}^{4}$ with $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, and $\mathbb{C}^{4}$ can be directly replaced by $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$ in that case. This means that the situation usually reported in modern manuals of quantum theory is exceptional and not the general one (Aerts & Sozzo,, 2014). In this general situation, where the marginal law conditions are empirically violated, no unique isomorphism exists between $\mathbb{C}^{4}$ and $\mathbb{C}^{2}\otimes\mathbb{C}^{2}$, hence entanglement cannot by captured only by the state. As a matter of fact, empirical violations of Equations (3) and (4) have been identified in cognitive domains, as anticipated above, but also in physical domains (Adenier & Khrennikov,, 2007, 2017; Bedhorz,, 2017; De Raedt Michielsen & Jin,, 2012; Kupczynski,, 2017). So far, however, relatively little attention in physics has been paid to the violation of the marginal law conditions in physical tests on entangled quantum quantities, and usually the explanation is sought in artifacts of the measurement process (Adenier & Khrennikov,, 2017). Because our research group has been one of the first to encounter empirical violations of the CHSH inequality in cognition and, at the same time, a violation of the marginal law conditions (Aerts & Sozzo,, 2011), our attention has been focused on the presence of both violations even in these early years, which includes situations where such violation appeared in tests on entangled physical entities (Adenier & Khrennikov,, 2007).
Thus, we devote the next section to outline the rationale underlying our approach to entanglement and its novelties with respect to more known approaches.
3 The novelties of our approach to entanglement
Let us outline how our research on the presence of entanglement in human cognition has influenced our general way of considering entanglement as a phenomenon, including how we consider the phenomenon in its physical presence (for a more detailed account of this, we refer to our various published articles on this topic, e.g., Aerts & Sozzo, (2014); Aerts et al., (2019)). Since the experimental situations we considered concern bipartite entities consisting of concepts that are combined and not physical entities that can be localized in space, as is the case for the study of entanglement in physics, we found it desirable to switch to a deeper characterization of the phenomenon concerning its non-classicality. More precisely, it should be a characterization that does not depend on the presence in space of the entities under consideration. This means that the notion usually associated with entanglement when the phenomenon occurs in physics, namely, ‘non-locality’, is not the notion desirably to use to characterize entanglement in the first place. More, even also the connection of the marginal law conditions with no-signaling should best be handled with caution in that respect, because there too the emphasis is on two entities that exist separately and, in some unspecified way, are each of them localized in space, often personificated by Alice and Bob, who can exchange signals with each other. The unconsciously powerful image thereby projected onto a situation of entanglement should not be underestimated in the already too narrow limitation it imposes on what entanglement really can be. Therefore, we prefer not to simply connect the situation in which the marginal law of Kolmogorovian probability is violated in an indiscriminate way with the physical imagined situation of no-signaling between Alice and Bob. We indeed have now become quite convinced, mainly through our research in quantum cognition, that entanglement confronts us with something ‘holistically deeper’, both when it occurs in human cognition and when it occurs in nature. Fortunately, there are plenty of notions that are indeed present in a well-defined way, and can be introduced mathematically rigorously in both physical and cognitive situations, that make it possible to characterize entanglement without taking recourse to an image which would stand in the way for a view that is closer to the nature of entanglement. Let us first bring out concretely the most important of these elements before returning to specifying how, in our opinion, we can also bring the image of entanglement closer to its deep nature.
In both physical and cognitive realms, the entities studied consist of two component entities. Although we have to immediately nuance this aspect, namely, that we do not know exactly in which way the two component entities are connected to each other and to the entity which is composed of them, finding out that ‘is’, by the way, the principle research aim on the phenomenon of entanglement, as originally pointed out already by Schrödinger. We have called such entities ‘bipartite entities’.
Probabilities are rigorously defined as limits of relative frequencies in both physical and cognitive realms, and this is where the first aspect of the characterization of entanglement comes in. There is entanglement if the probabilities obtained from the joint measurements on the bipartite entity are not a factorization of the probabilities obtained from the separate measurements that make up the joint measurement, the bipartite entity being prepared in a well-defined state. Thus, this means that we consider the notions of ‘joint measurement’ and ‘bipartite entity prepared in a well-defined state’ to be elements needed to characterize entanglement, and indeed, for the cognitive situations we studied, we have always taken care of these elements to be well-defined.
What becomes the ‘non-classicality’ of this general characterization of entanglement? Well, if classical physical entities are separated from each other in space, then the aforementioned probabilities factorize. And this archetypical situation that we know well, even from our everyday experience with classical entities, is why notions of non-locality and no-signaling made their appearance in the study of entanglement. Also, when bipartite quantum entities described by the tensor product procedure are in a product state, and the joint measurements are represented by product operators, then the said probabilities factorize. That means that in a situation modelled with the quantum formalism, if these probabilities do not factorize, at least one of these aspect fails. The most studied failing aspect is when the state of the bipartite entity is not a product state, but a superposition of product states, which is customarily now called an entangled state. Little attention was paid to the situation where the joint measurements are not represented by product operators, which also leads to a situation where the probabilities do not factorize. In what follows we will analyze why this is the case and why our approach therein differs from the most common one.
What is the status of the violation of the CHSH inequality within the more general framework in which we want to study both physical and cognitive situations with entanglement? Here we use Pitowsky’s study of correlation polytopes and their connection with classical Kolmogorovian probabilities (Pitowsky,, 1989). This theory proves that satisfying the CHSH inequality in all its variants is equivalent to the existence of a classical Kolmogorovian probability model for the corresponding probabilities. Pytowsky’s theory thus shows that the violation of one of the CHSH inequalities is sufficient for the presence of a non-classical probability structure. This does not mean that it must necessarily be a quantum probability structure, but it does justify the motivation to try quantum probability as a modelling tool, which is what we did in our research.
Little attention has been paid to the situation where entanglement has its origin in violating both the CHSH inequality and the marginal law conditions. The reason is that violating the marginal law conditions which is equivalent to violating the no-signaling conditions is by many interpreted as indicating the presence of signaling, and therefore these situations are misidentified by many physicists as situations where signaling is present, and thus trivially uninteresting in terms of the phenomenon of entanglement. What is the error in the reasoning made here? The no-signaling conditions only constitute a ‘sufficient’ condition for no-signaling to be present, but it is not a ‘necessary’ condition. From a logical perspective, this is a rather subtle situation where typically logical reasoning errors are made, and we believe this is frequently the case here. Thus, we stress that, even if the no-signaling conditions are not satisfied, it is perfectly possible that no signals (can) be sent between Alice and Bob.
Let us now bring out the elements that motivate us to interpret the violation of the marginal law conditions as being due to a lack of presence of some form of symmetry in the considered situation rather than as a consequence of the existence of signaling.
Already in the early days of studying entanglement, around the same period of Aspect’s experiments, one of the authors of this article proposed an explicit mechanism that is abundantly present in nature (and later identified in human cognition) in a subtle but obvious way, and that violates Bell’s inequalities (Aerts,, 1982, 1983). The mechanism starts from the mere potential presence of correlations in the state of the bipartite entity – in the example worked out in Aerts, (1982, 1983), this potential presence is the state of water contained in two vessels connected by a tube, and that potential is actualized by and during the joint measurements themselves – in the example in Aerts, (1983), the water volumes extracted from each vessel by siphons will end up correlated when collected in outcome reference vessels due to the connecting tube and the classical principle of communicating vessels (see Figure 1). The group of scientists studying the phenomenon of entanglement in those early days was tiny and without the authority to deal with anything important. As a consequence, all sorts of hypotheses were considered. Even the hypothesis that quantum theory may have wrongly predicted a violation of Bell’s inequalities was among the opinions of quite a few members of that little group, including Bell himself, but also Clauser’s motivation was to find where quantum theory was wrong. Things turned out differently and entanglement showed finally to be a phenomenon real in all its aspects predicted by quantum theory. The common way of looking at entanglement in those days, expressed by Einstein as ‘spooky action at-a-distance’, seemed to carry an irreconcilable conflict with relativity theory. Meanwhile, however, the no-signaling conditions and their structural equivalence with the marginal law conditions also became known, hence entanglement could not be used to send signals between Alice and Bob.
Figure 1: A schematic representation of the vessels of water example worked out in Aerts, (1982, 1983) (see also the discussion in Aerts & Sassoli de Bianchi, (2021)).
The example of a violation of Bell’s inequalities from the presence of potential correlations only actualizing pduring the joint measurements proposed in Aerts, (1982, 1983) also violates the no-signaling conditions. Therefore, later on, when in the 1990s a much larger group of physicists and computer scientists started analyzing quantum structures, mainly out of interest in quantum computation, and somewhat later also quantum information, it was classified as a mechanism that is not really interesting, since identified as violating Bell’s inequalities because of containing signaling. However, However, later, variants of the same mechanism were created where the no-signaling conditions are not violated (Aerts,, 1991, 2005) and, inspired by this, an example was also worked out in human cognition, with a combination of concepts, where the same mechanism causes a violation of Bell’s inequalities while satisfying the no-signaling conditions (Aerts,, 2014). If these examples are compared, considering the proposed mechanism in detail, it is clear that whether or not the no-signaling conditions are satisfied is mainly a matter of presence of enough symmetry in the mechanism, but that the mechanism, even in the cases where due to absence of the necessary symmetry the no-signaling conditions are violated, contains no-signaling nevertheless. We hereby recall again that the no-signaling conditions only constitute a sufficient condition in connection with no-signaling but not a necessary condition, as it is often considered erroneously to be.
A second element related to our interpretation of the violation of the marginal law conditions in most experiments in human cognition that violate Bell’s inequalities, as being due to a ‘lack of symmetry’ and not to the presence of signaling, consists in a reappraisal of the mechanism itself. More than before, we have become convinced that the mechanism proposed in the various examples Aerts, (1982, 1983, 1991, 2005, 2014), is also the mechanism that violates Bell’s inequalities in physics and let us make the analysis for the case of the singlet spin state that supports us in this conviction. The singlet spin state is a superposition of two product states, associated with the pairs (up, down) and (down, up) of the respective spins of the two component entities of the composite bipartite entity, where the (up, down) and (down, up) can be in any possible space direction and its opposite. That means it is a state that only potentially holds correlations, where these correlations are actualized during the spin measurements. This is exactly the mechanism by which the examples above were constructed, and the macroscopic model in Aerts, (1991) realizes the exact probabilities and states within two coupled Bloch spheres. By the way, we can also understand after this analysis why no spooky action at-a-distance takes place. The collapse as a result of the measurement happens indeed from the singlet spin state as superposition to one of those product states, (up, down) or (down, up) in some space direction, even if meanwhile the entities carrying the spin are far away from each other in space. The collapse is in that sense a change of state which, however, is not at all connected with whether or not the entities are far apart in space, and it gives rise to outcome events that are space-like separated. Also for the examples elaborated in Aerts, (1982, 1983, 2005, 2014), the distance between the outcome regions does not play any role in the dynamics that plays out for the outcome to occur and the outcomes are also events that are space-like separated. The examples in Aerts, (1982, 1983, 1991, 2005, 2014) are macroscopic and do not contain microscopic quantum elements, does this mean that we believe entanglement also admits a macroscopic variant? Our answer to this question is positive with the nuance that, as is the case for microscopic realizations of entanglement, a correlation coherence region comes into play, where it is the quality of the experimental setup that should ensure that this correlation coherence region is preserved during the experimental test and is large enough to contain the outcome regions. It should not be surprising that in the macroscopic realm with rather classical entities, it is possible to realize a ‘potentiality for correlations’, think of how potential energy in the same classical environment realizes a ‘potentiality for energy’. But also note that for the macroscopic physical examples it is necessary that there remains a connection in space between the two part entities which is not necessary for the quantum microscopic examples, and one could say that this is the way that non-locality manifests itself in our approach. We can see for this aspect the connection with a recent perspective on entanglement that we developed, where it may have its origin in human cooperation to combat uncertainty, not occurring spontaneously in that sense in the macroscopic world, which may be a difference from the more spontaneous way it occurs in the microscopic quantum world (Aerts et al., 2023a, ).
There remains one more element that needs explanation, namely, how it is that the size of the CHSH inequality expression for cognitive tests, and likewise for the macroscopic examples, can exceed Cirel’son’s bound, while theoretically this is not the case for microscopic quantum entities that are entangled. We can understand the condition for the existence of this bound when we consider Cirel’son’s proof of it in detail. We can then note that it is necessary to be able to represent the four considered joint measurements present in the CHSH inequality expression by one self-adjoint operator in the considered Hilbert space. This indicates the supposed presence of a very large and rigorous quantum coherence that incorporates the four joint measurements simultaneously and independently of how and when they may be performed separately. This very large internal coherence is not easy to bring present in the cognitive tests, nor in the macroscopic tests, although it was realized in Aerts, (1991), which gives rise to identical probabilities as the spin of a spin 1/2 quantum entity. The other examples violate Cirel’son’s bound, and this is probably linked to the same lack of symmetry present, which in most cognitive tests, and in the most obvious macroscopic examples, also violates the marginal law conditions.
It is already mentioned that our investigation of the phenomenon of entanglement in quantum cognition made us realize that something of a deeper holistic nature is going on when two entities enter into a situation of entanglement with each other. We have now brought out all the elements to make this also explicit. The entangled entity is more than usually supposed to be a new entity in itself only still connected to the original entities from which it was formed in a very specific way, and this has become even more clear to us by analyzing and understanding more and more deeply how concepts are entangled in human language. We recently introduced a new expression ‘contextual updating’ taking place in human language relative to the global meaning carried by the whole context, which contains an important part of our new understanding. That also the marginal law conditions are violated as a consequence of the entangled entity behaving fully as a new entity relative to the global meaning context is clearly identified in language, and it is also clearly seen that this has nothing to do with the presence of signals. We believe that a similar process takes place when two physical entities transform into an entangled entity, namely, a new entity is created that forms itself not primarily with respect to the two component entities, but contextually with respect to the global quantum coherence present, in which the two component entities will generally play an important role, but in principle not only they (Aerts et al., 2023a, ) (see Section 6). In this sense, we think that the situation where both the CHSH inequality and the marginal law conditions, are violated is the default situation in terms of entanglement, namely, the situation where its intrinsic reality is best expressed. Admittedly, it remains important to verify that there is no-signalling present that causes the violations, if one wants to distinguish the cases of non-signaling entanglement from the cases where the entanglement is provoked by signaling. We remind however that, even when signaling is involved, a non-classical probability is generated by it, and the phenomenon remains non-classical.
@@ 3 The novelties of our approach to entanglement
The considerations above led us to investigate in detail the above mentioned identification problem and to elaborate a quantum mathematical representation of the Bell-type tests that violate both the marginal law conditions, the CHSH inequality and Cirel’son’s bound. In this quantum theoretical framework, we explicitly introduce entangled it measurements (Aerts & Sozzo,, 2014; Aerts Arguëlles & Sozzo,, 2020; Aerts et al.,, 2021; Aerts et al., 2023b, ). Specifically, we have proved that, whenever the concepts Animal and Acts combine to form the combination The Animal Acts, a strong form of entanglement is created between Animal and Acts, which is such that, not only the state of the composite entity The Animal Acts is entangled, but also the joint measurements are entangled.
@@
We will see in Sections 5 and 6 that the entanglement above is due to the peculiar way in which the meaning of The Animal Acts relates to the meanings of Animal and Acts, which violates the classical semantic rules of composition. Before doing this, however, we need to present the details of the novel cognitive test based on videos which we have recently performed on this conceptual combination. This will be the aim of Section 4.
@@ 4 A novel video-based cognitive test
A sample of 221 individuals were presented in a ‘within subjects design’ a HTML5 questionnaire which contained four joint measurements $AB$, $AB^{\prime}$, $A^{\prime}B$ and $A^{\prime}B^{\prime}$ whose setting was similar to the typical setting of a Bell-type test presented in Section 2. More specifically, participants were preliminarily asked to read an ‘introductory text’ where an explanation of the type of judgement test they had to complete and a description of the tasks involved in the judgement test were provided. More specifically, the participants had to preliminary complete a simple introductory test on the concept Fruit, judging the item that they considered as a good example of the concept Fruit. The items of Fruit, indicated in Figure 2, were generated by image-based generative intelligence, and we proposed the item Strawberry, which is an enlarged portion of an inflorescence, and Tomato, which is a fruit, to draw maximum attention to the choice. Cherry and Pomegranate were the remaining items of Fruit to judge upon.
@@
Then, in each joint measurement, participants were asked to choose which item in a list of four items they judged as a good example of the conceptual combination The Animal Acts. In each joint measurement, each of the four possible items of The Animal Act was a video showing an animal and the corresponding action. More specifically, the videos, indicated in Figure 3, were created as follows:
@@
Figure -1: sub-test 1
Figure 0: sub-test 2
@@
Before analysing the results of the test, it is worth to make some remarks on the choice of its setting. The detection of entanglement in Bell-type cognitive tests, indeed, relies on a precise preparation of both the initial state and the joint measurements. Insufficient initial conditions, as lacking a brief introduction and prior knowledge of the animals and sounds involved, can hinder the ability to identify entanglement. Birds could offer an interesting case study because some species possess unique vocalizations distinct from other individuals. This individuality makes them well-suited for entanglement experiments involving experts. One could, e.g., consider, two pairs of items of Animal, namely, (Cuckoo, Nightingale) and (Thrush, Goldfinch), and two pairs of items of Acts, (Cuckoos, Trills) and (Whistles, Snails), without determining a violation of the CHSH inequality in Equation (1). In other words, cognitive tests on unskilled participants may fail to detect entanglement or yield lower value of the CHSH factor $\Delta_{CHSH}$ in Equation (2). A similar situation is likely to occur if the actions are not typical of the animals. Consider, e.g., two pairs of items of Animal (Horse, Bear) and (Dog, Camel), and two pairs of items of Sport (Racing, Fishing) and (Sledding, Show Jumping) for the conceptual combination The Animal does Sport. Also in this case, we expect the CHSH inequality in Equation (1) not to be violated or to show a little violation. This is why the identification of the items to use in the cognitive test followed strict prescriptions, as follows:
@@
In the joint measurement $AB$, participants had to choose the best example of The Animal Acts within the four items:
@@
In the joint measurement $AB^{\prime}$, participants had to choose the best example of The Animal Acts within the four items:
@@
In the joint measurement $A^{\prime}B$, participants had to choose the best example of The Animal Acts within the four items:
@@
Finally, in the joint measurement $A^{\prime}B^{\prime}$, participants had to choose the best example of The Animal Acts within the four items:
@@
For each joint measurement $XY$, we collected the relative frequencies of the obtained responses which we considered, in the large number limit, as the probability $\mu(X_{i}Y_{j})$ that the outcome $X_{i}Y_{j}=\pm 1$ is obtained in the corresponding measurement, $X=A,A^{\prime}$, $Y=B,B^{\prime}$. Table 1 reports the judgement probabilities computed in this way. Referring to these probabilities, we can then calculate the expectation values, or correlation functions, of the joint measurements $AB$, $AB^{\prime}$, $A^{\prime}B$ and $A^{\prime}B^{\prime}$, as follows:
@@
The numerical value $2.9774$ exceeds the classical limit imposed by the CHSH inequality in Equation (1) and is also above Cirel’son’s bound $2\sqrt{2}\approx 2.8284$. A simple check in Table 1 reveals that the marginal law conditions in Equations (3) and (4) is also systematically violated here. This empirical pattern confirms and strengthens the results obtained in previous text-based cognitive tests on the conceptual combination The Animal Acts, namely Aerts & Sozzo, (2011); Aerts et al., 2023b ; Bertini Lemporini & Moriani, (2023). Table 2 reports the judgement probabilities and the CHSH factor in these three text-based tests.
By comparing the CHSH factors in Tables 1 and 2, we can see that a systematic violation of the CHSH inequality occurs in all tests. We also notice that the value of the CHSH factor in the test in Aerts & Sozzo, (2011) is far from Cirel’son’s bound, the test in Aerts et al., 2023b is close to that bound, and the test in Bertini Lemporini & Moriani, (2023), together with the present test violate Cirel’son’s bound. That the test in Aerts & Sozzo, (2011) has a relatively lower deviation from classicality than the other tests might be due to the fact that the introductory text in Aerts & Sozzo, (2011) encouraged the participants to also take into account emotions and imagination in their judgements. This led them to make choices that are less natural, as The Bear Meows, The Cat Growls, or The Tiger Meows, thus determining a relatively lower violation of the CHSH inequality in Equation (1).